Hello!
How much water should be added to 1 gallon of pure antifreeeze to obtain a solution that is 60% antifreeze?
If someone could show me how to set this problem up it would be appreicated.

Answer by scott8148(6628)

1(100%)+x(0%)=(1+x)(60%) ... 1(1)+x(0)=(1+x)(.6) ... 1=.6+.6x ... .4=.6x ... 2/3=x

pure antifreeze is 100% antifreeze, water is 0% antifreeze